# Formatting Instructions for ICLR 2025 Conference Submissions

### v1

@@ 4.1 Application 1: Depth Pruning
As a first application, we apply EvoPress on Depth Pruning. Although removing entire transformer blocks generally results in greater accuracy losses compared to other compression techniques, this approach recently attracted attention in the context of initializing smaller models, as it guarantees speedups proportional to the sparsity (Samragh et al., 2023; Kim et al., 2024). Additionally, block dropping provides insights into the capabilities of transformer models, making it relevant for interpretability. We will compare against the following baselines:
  • Shortened Llama (Kim et al., 2024): Scores blocks on the perplexity change after removal.
  • ShortGPT (Men et al., 2024): Blocks are scored based on the average cosine similarity between input and output embeddings, including the residual stream.
  • Weight Subcloning (Samragh et al., 2023): Blocks are scored using the ratio $||f(x)||/||f(x)+x||$, where $x$ is the input embedding and $f(x)$ is the block’s output, excluding the residual stream.
  • Sliding Window Cosine Similarity (Gromov et al., 2024): Sets of consecutive blocks are scored based on the cosine similarity between embeddings before and after the blocks, including the residual stream.
While Gromov et al. (2024) directly score entire removal configurations, Shortened Llama, ShortGPT, and Weight Subcloning determine block removals based on their isolated scores.
Figure 2: Depth pruning results, on Mistral-7B-v0.3. (Left) Relative to all prior methods, EvoPress shows significantly lower PPL gap relative to the uncompressed model, with remarkably large gaps at medium compression rates. (Right) Examining the blocks dropped, we observe that EvoPress isolates completely different profiles relative to ShortGPT (which scores by cosine similarity).
@@
Experimental Results.
Figure 2 compares our method with baselines from previous work on Mistral-7B-v0.3. For a better comparison, we also included results where only entire transformer blocks are removed (Attn.+MLP). EvoPress consistently outperforms all previous methods, showing significant improvements even at medium sparsity levels. While all baseline methods fail entirely beyond 31.25% sparsity, EvoPress identifies functional submodels even when removing half of the model. To our knowledge, this is the first method to achieve such results. We observed similar collapses in Llama-2-7B, Llama-3-8B and Llama-3.1-8B. Overall, EvoPress consistently outperforms all baselines across all tested models and sparsities (see Appendix D.1 for full results), and does so in a matter of minutes (Appendix D.2).
All four previous methods rely on human-crafted scoring methods to identify the optimal combination of transformer blocks to remove. However, these approaches are not only suboptimal, but also prone to bias, as their results may reflect the characteristics of the method itself rather than the model’s true behavior. Specifically, we found that most scoring methods tend to favor deeper blocks, resulting in highly similar removal configurations across different prior scoring methods (Appendix Table 13). This likely occurs because methods that bias towards deeper blocks generally perform better than those that focus on earlier blocks, although neither may be optimal. In contrast, EvoPress employs an unbiased approach, offering more accurate and meaningful insights into the model. As shown in Figure 2, we found that the deeper layers are not necessarily the least important, contradicting conclusions drawn in prior work (Gromov et al., 2024; Men et al., 2024).
@@ 4.2 Application 2: Unstructured Sparsity
Sparsity levels are generated as follows: For each layer, we first produce the base level corresponding to the targeted average sparsity. Then, we generate both higher and lower compression levels, where the difference between two levels corresponds to a fixed number of weights. In our experiments, we used a “step size” of 1M weights uniformly. This approach enables the mutation of compression levels across all layers, independently of their size. We adopt SparseGPT (Frantar & Alistarh, 2023) as a fast and efficient one-shot layer pruner.
@@
[…]
Mistral-7B-v0.3 Dense 4.82 7.72 48.9 79.6 60.9 80.3 73.9 68.7
Uniform 23.08 30.03 27.1 60.9 36.1 65.9 59.4 49.9
OWL 17.22 21.66 27.9 62.6 38.6 67.0 63.5 51.9
EvoPress 14.42 16.46 31.6 64.7 41.4 69.5 61.9 53.8
Llama-3-8B Dense 5.54 7.10 50.4 80.1 60.2 79.7 72.6 68.6
Uniform 85.84 98.35 22.7 49.9 31.4 62.1 54.4 44.1
OWL 48.07 52.32 27.0 54.9 36.6 65.1 58.6 48.4
EvoPress 28.76 33.72 28.9 56.7 38.6 68.0 61.7 50.8
Llama-3.1-8B Dense 5.61 8.90 51.2 81.4 60.0 80.1 73.9 69.3
Uniform 68.97 103.27 22.3 49.7 31.5 61.6 55.6 44.2
OWL 42.29 48.65 27.4 55.8 36.5 65.7 60.7 49.2
EvoPress 24.32 30.58 29.1 62.4 39.5 68.9 60.3 52.0
Phi-3-Medium-14B Dense 4.02 8.31 60.9 84.1 64.0 81.0 76.2 73.2
Uniform 16.66 24.73 36.9 70.6 40.0 69.4 65.8 56.5
OWL 15.66 23.38 35.7 69.2 39.4 68.3 64.4 55.4
EvoPress 13.83 19.13 41.5 73.0 43.6 71.8 69.1 59.8
@@
Examining sparsity profiles (Appendix Figures 8 and 9), we observe that EvoPress prunes the first blocks less aggressively, blocks in the beginning of the second half of the model more aggressively while keeping the even deeper blocks relatively dense. Notably, EvoPress assigns high importance to the v_proj matrices, reducing its sparsity to below 45%, compared to an overall average of 70%.
@@
EvoPress is also time-efficient. Figure 3 illustrates the rapid convergence of our method vs. iterations and time, with smooth and steady improvements in test perplexity. Moreover, we found that, by significantly reducing the number of tokens used in the multi-step selection evaluation, by 4×4$\times$4 × in the first step and 8×8$\times$8 × in the last step, and making each generation have fewer offspring, we can significantly speed up the search. This “super-fast” version converges in a little over one GPU hour to similar test PPL (Figure 3, right), demonstrating the sample-robustness of EvoPress, which can lead to further efficiency gains.
Figure 3: Left: The convergence of EvoPress vs. number of generations and wall-clock time (on a single RTX 3090 GPU with 24GB RAM) for Llama-2-7B. We observe convergence close to optimum in 5-6h; Right: Convergence of the “super-fast” version which reduces the number of tokens used for each evaluation. It converges to similar accuracy in little over one hour, in the same setting.
@@ 4.3 Application 3: Quantization
Finally, we apply EvoPress to the more challenging task of non-uniform neural network quantization, where the widely adopted baseline is uniform per-layer quantization (Frantar et al., 2022; Lin et al., 2023; Chee et al., 2023). However, one could expect that different layers exhibit varying sensitivity to quantization, similar to the behavior observed with unstructured sparsity. Furthermore, most quantization techniques are restricted to certain bitwdiths; for example, scalar quantization methods like GPTQ are limited to integral bitwidths, with only a few of these bitwidths being practical for current hardware implementations. Therefore, non-uniform quantization is critical for achieving a broader range of compression targets.
As baselines, we consider uniform quantization and DP search. (While OWL has also been applied to quantization, the authors found that it underperforms even relative to uniform per-layer quantization (Yin et al., 2024).) The DP search is very similar to SPDY (Frantar & Alistarh, 2022), where the goal is to minimize the Normalized Mean Squared Error (NMSE), defined as $\mathrm{NMSE}=\|\hat{Y}-Y\|_{2}^{2}/\|Y\|_{2}^{2}$, where $Y$ represents the original model output at a layer, and $\hat{Y}$ the output of the compressed model. Then, the optimal compression allocation can be determined via a dynamic programming (DP) approach. The full SPDY method applies a second iterative random search step, which is very expensive to implement at LLM scale, and is therefore omitted. We create configurations with varying bitwidths and run EvoPress to determine the optimal configuration for the target compression ratio.
@@
[…]
Mistral-7B-v0.3 Dense 4.82 7.72 48.9 79.6 60.9 80.3 73.9 68.7
Uniform 5.54 8.57 47.0 77.3 58.6 78.5 70.0 66.3
DP 5.79 8.84 46.3 77.3 57.5 79.33 69.2 66.0
EvoPress 5.21 8.42 48.2 78.5 58.9 79.2 71.0 67.1
Llama-3-8B Dense 5.54 7.10 50.4 80.1 60.2 79.7 72.6 68.6
Uniform 12.19 15.76 35.2 66.9 54.0 75.2 69.6 60.2
DP 29.00 20.03 39.8 72.0 52.9 74.7 67.2 61.3
EvoPress 7.49 12.03 43.0 76.4 55.4 77.3 69.7 64.3
Llama-3.1-8B Dense 5.61 8.90 51.2 81.4 60.0 80.1 73.9 69.3
Uniform 27.33 12.44 39.3 71.8 53.2 76.3 70.6 62.2
DP 36.66 15.63 28.0 56.0 51.6 68.9 68.0 54.5
EvoPress 7.07 11.90 42.2 75.0 55.9 76.9 71.0 64.2
Phi-3-Medium-14B Dense 4.02 8.31 60.9 84.1 64.0 81.0 76.2 73.2
Uniform 5.18 9.05 55.1 81.6 60.8 78.9 73.6 70.0
DP 5.72 9.71 54.7 80.4 58.4 78.6 73.5 69.1
EvoPress 5.09 9.00 56.7 82.6 61.0 79.2 74.7 70.8
@@
To validate the effectiveness of EvoPress, we consider the challenging problem of quantization to 3 bits and below. For this compression rate, uniform GPTQ quantization faces significant performance drops, motivating a dynamic quantization bitwidth allocation. We produce quantization levels at 2, 3, 4, 5, and 6 bits and search for an optimal compression profile with respect to the fitness function. The results in Table 3 indicate that non-uniform quantization with EvoPress produces superior models compared to the baseline methods. The improvements are even more pronounced at 2.25 bit and 2.5 bit quantization, as detailed in Appendix F.1.
We visualize quantization configurations found by EvoPress for Llama-3.1-8B in Appendix Figures 11 and 12. Specifically, we observe that the second and final blocks are compressed less aggressively, while the first block undergoes the highest level of compression. This contrasts with unstructured sparsity, where the first block is among the least compressed. These findings highlight that dynamic compression allocation must account for the specific compression method used, providing insight into OWL’s failure to generalize and underscoring the need for automated compression allocation through EvoPress.
Overall, we observe that, in this case as well, EvoPress yields significant accuracy improvements (e.g., 4.1 and 2.0 points on the zero-shot averages for Llama-3-8B and Llama-3.1-8B, respectively), compared to the uniform profile. Moreover, the improvement over the next-best method is always significant, both in terms of perplexity and zero-shot accuracy.
@@ 5 Conclusion
We have presented EvoPress, a unified optimization framework for non-uniform compression. EvoPress is based on a new provably-convergent evolutionary search algorithm with low sample and iteration complexity, that is especially well-suited to the loss landscapes arising in LLM compression. Specifically, we have shown that EvoPress can converge extremely fast to accurate configurations for various non-uniform LLM compression problems, and is also fast to execute in practice. We also emphasize the breadth of our study, our method was implemented and tested on three different compression approaches, relative to prior work which largely focused on a single application. Experimental results showed that EvoPress consistently outperforms prior dynamic compression approaches, across all compression types, with large gaps at medium to high compression.
Limitations.
One interesting direction we did not investigate is the possibility of combining different compression approaches into the same search space. This would require changes to our switch mutation strategy, but should be feasible in most cases. Second, we did not investigate finer-grained structured pruning (i.e., removing rows and columns from the weight matrices), as it usually requires extensive retraining to recover accuracy. We plan to investigate this in future work. Finally, we plan to extend our quantization results, to show end-to-end speedups in the context of an inference engine supporting multiple compressed formats, such as vLLM (Kwon et al., 2023).
@@ A.1 A Warm-Up Argument for a Single Offspring
The overall goal of this section is to prove Theorem 1. As the main argument is quite complex, relying heavily on stochastic drift analysis, we begin with a warm-up, namely by presenting a simpler proof for the restricted case where $\lambda=1$.
@@
Theorem 2 (Single offspring).
@@ Discussion.
It is important to note that the above proof relies on the key assumption that whenever one of the $j^{2}$ ”good” pairs is selected during mutation, the resulting offspring is the fittest among all candidates. This condition holds naturally when there is only a single offspring, as the offspring produced by flipping one of the $j^{2}$ pairs will have higher fitness than the parent. However, in the case of multiple offspring, this approach breaks down, as an offspring produced by flipping one of the $j^{2}$ ”good” pairs might still have lower fitness than another offspring that was not generated by flipping one of these $j^{2}$ ”good” pairs.
@@ A.2 The Main Argument
Drift analysis, originally developed to study all kinds of random walks and Markov chains, has become the most widely used technique for analyzing the runtime of evolutionary algorithms in recent years. It works by first defining a potential function $X^{(t)}$ that measures the progress over each step $t$ of the optimization. By estimating how this potential changes at each step in expectation, i.e., computing the drift in $X^{(t)}$, one can then make probabilistic statements about the number of steps required until the potential reaches a certain threshold, also called the hitting time. To this end, a variety of drift theorems have been established, two of which will be employed in our proof. For a more thorough introduction to drift analysis, we refer to Lengler (2020).
@@
Theorem 3 (Multiplicative Drift, Tail Bound (Doerr & Goldberg, 2010)).
@@
Theorem 4 (Variable Drift Theorem (Johannsen, 2010)).
@@
As in the proof of Theorem 2 let $w\in\mathbb{R}^{n}$ represent the weights associated with a linear function of the form $f(x)=\sum_{i=1}^{n}x_{i}\cdot w_{i}$. To establish an upper bound, we can again assume that no two weights are equal. Additionally, without loss of generality, assume that the weights are ordered in increasing value, i.e., $w_{1}<w_{2}<\dots<w_{n}$, and that $k\leq n-k$, as the other case follows by symmetry. Let $x^{(t)}$ denote the search point at step $t$, and define
@@
which captures the number of inversions at step $t$. Since $x_{\text{opt}}=0^{k}1^{n-k}$ is the only bit string with $k$ 00-bits without inversions, we have $X^{(t)}=0$ if and only if $x^{(t)}=x_{\text{opt}}$. At the same time, no bit string with $k$ 00-bits has more than $k(n-k)$ inversions, hence, $X^{(t)}\leq k(n-k)$ at all times. During mutation, each of the $\lambda$ offspring is generated independently by copying the parent $x^{(t)}$, choosing uniformly at random one of the $1$-bits, choosing uniformly at random one of the 00-bits and finally flipping both bits. This flipping can also be viewed as switching both bits, so that bits ”move” across the search point in consecutive generations. We will use this abstraction in a later step of the proof.
@@
At the same time, switching two bits corresponding to an inversion decreases the number of inversions by the difference in their positions, which we call the spread of an inversion. This implies that any offspring fitter than its parent must have fewer inversions than its parent and therefore, $X^{(t+1)}\leq X^{(t)}$ for all $t$. Note that we cannot make the same statement about the entire group of offspring, meaning, the fittest offspring is not guaranteed to have the fewest inversions. Since $X^{(t)}$ is non-increasing we can decompose $T$ into the number of steps required until for the first time the current search point $x^{(t)}$ has at most $\frac{k(n-k)}{\lambda}$ inversions and the number of steps required from there until the optimum is found. By linearity of expectation
@@
Denote by $\hat{T_{j}}$ the number of steps required until $Z_{j}$ reaches $1$, only counting steps where $Z_{j}$ is decreased. Using a tail bound for the Multiplicative Drift Theorem (Theorem 3) we have that
@@
Finally, as the probability of having less than $Ck(n-k)/\lambda$ ”successful” generations in the considered time period is sub-constant, we conclude via another union bound that there exists a constant $C^{\prime}$ such that event $B_{C}$ occurs with probability at least $1/C^{\prime}$. Consequently, we have
@@
Lemma 1 indicates that when selecting an offspring uniformly at random from all those with higher fitness than the parent (i.e., those generated by flipping an inversion), the expected number of inversions in that offspring is at least $\sqrt{s}/16$ fewer than in the parent. We can now reveal the randomness in two steps. First, we only uncover how many of the generated offspring are fitter than the parent. Given that there is only a single fitter offspring, i.e., conditioned on $A_{1}$, we then uncover its number of inversions. Clearly, this single fitter offspring is now sampled uniformly at random from all offspring with higher fitness than $x^{(t)}$; thus, for $s\leq k(n-k)/\lambda$
@@
Finally, applying Johannsen’s Variable Drift Theorem (Johannsen, 2010) (Theorem 4) yields

### v2

@@ 4.1 Application 1: Depth Pruning
We first apply EvoPress on Depth Pruning. Although removing entire transformer blocks generally results in large accuracy losses, this approach recently attracted attention in the context of initializing smaller models, as it guarantees speedups proportional to the sparsity (Samragh et al., 2023; Kim et al., 2024). Additionally, block dropping provides insights into the capabilities of transformer models, making it relevant for interpretability. We will compare against the following baselines: (1) Shortened Llama (Kim et al., 2024), which scores blocks on the perplexity change after removal; (2) ShortGPT (Men et al., 2024), where blocks are scored based on the average cosine similarity between input and output embeddings, including the residual stream; (3) Weight Subcloning (Samragh et al., 2023), where blocks are scored using the ratio $||f(x)||/||f(x)+x||$, where $x$ is the input embedding and $f(x)$ is the block’s output, excluding the residual stream; (4) Sliding Window Cosine Similarity (Gromov et al., 2024), where sets of consecutive blocks are scored based on the cosine similarity between embeddings before and after the blocks, including the residual stream. While Gromov et al. (2024) directly score entire removal configurations, the other approaches determine block removals based on their isolated scores.
Figure 2: Depth pruning results, on Mistral-7B-v0.3. Relative to all prior methods, EvoPress shows significantly lower PPL gap relative to the uncompressed model, with remarkably large gaps at medium compression rates.
Figure 3: Left: The convergence of EvoPress vs. number of generations and wall-clock time (on a single RTX 3090 GPU with 24GB RAM) for Llama-2-7B. We observe convergence close to optimum in 5-6h; Right: Convergence of the “super-fast” version which reduces the number of tokens used for each evaluation. It converges to similar accuracy in little over one hour, in the same setting. The KL-Divergence corresponds to the fitness of the survivor in each generation, which is measured on a random minibatch of the entire training dataset. The perplexity is computed on the entire test dataset.
@@
Results.
Figure 2 compares our method with baselines from previous work on Mistral-7B-v0.3. For a better comparison, we also included results where only entire transformer blocks are removed (Attn.+MLP). EvoPress consistently outperforms all previous methods, showing significant improvements even at medium sparsity levels. While all baseline methods fail entirely beyond 31.25% sparsity, EvoPress identifies functional submodels even when removing half of the model. To our knowledge, this is the first method to achieve such results. We observed similar collapses in Llama-2-7B, Llama-3-8B and Llama-3.1-8B. Overall, EvoPress consistently outperforms all baselines across all tested models and sparsities (see Appendix D.1 for full results), and does so in a matter of minutes (Appendix D.2). We provide runtime comparisons as well as additional comparisons to the iterative search methods SLEB (Song et al., 2024) and BlockPruner (Zhong et al., 2024) in Appendix G.
All four previous methods rely on human-crafted scoring methods to identify the optimal combination of transformer blocks to remove. This is not only suboptimal, but also prone to bias, as their results may reflect the characteristics of the method itself rather than the model’s true behavior. Specifically, we found that most scoring methods tend to favor deeper blocks, resulting in highly similar removal configurations across different prior scoring methods (Appendix Table 13). This likely occurs because methods that bias towards deeper blocks generally perform better than those that focus on earlier blocks, although neither may be optimal. In contrast, EvoPress employs an unbiased approach, offering more accurate and meaningful insights into the model.
@@ 4.2 Application 2: Unstructured Sparsity
Sparsity levels are generated as follows: For each layer, we first produce the base level corresponding to the targeted average sparsity. Then, we generate both higher and lower compression levels, where the difference between two levels corresponds to a fixed number of weights. In our experiments, we used a “step size” of 1M weights uniformly. This approach enables the mutation of compression levels across all layers, independently of their size. We adopt SparseGPT (Frantar & Alistarh, 2023) for layer pruning. We provide results of EvoPress with Wanda pruning (Sun et al., 2024) in Appendix G.
@@
Model Method Wiki2$\downarrow$ C4$\downarrow$ Task Avg.$\uparrow$
Mistral-7B-v0.3 Dense 4.82 7.72 68.7
Uniform 23.08 30.03 49.9
OWL 17.22 21.66 51.9
EvoPress 14.42 16.46 53.8
Llama-3-8B Dense 5.54 7.10 68.6
Uniform 85.84 98.35 44.1
OWL 48.07 52.32 48.4
EvoPress 28.76 33.72 50.8
Phi-3-Medium-14B Dense 4.02 8.31 73.2
Uniform 16.66 24.73 56.5
OWL 15.66 23.38 55.4
EvoPress 13.83 19.13 59.8
@@
Examining sparsity profiles (Appendix Figures 12 and 13), we observe that EvoPress prunes the first blocks less aggressively, blocks in the beginning of the second half of the model more aggressively while keeping the even deeper blocks relatively dense. Notably, EvoPress assigns high importance to the v_proj matrices, reducing its sparsity to below 45%, compared to an overall average of 70%.
@@
EvoPress is also time-efficient. Figure 3 illustrates the rapid convergence of our method vs. iterations and time, with steady improvements in test perplexity. Moreover, by reducing the number of tokens used in the multi-step selection evaluation, by 4×4$\times$4 × in the first step and 8×8$\times$8 × in the last step, and making each generation have fewer offspring, we can significantly speed up the search. This “super-fast” version converges in a little over one GPU hour to similar test PPL (Figure 3, right), demonstrating the robustness of EvoPress, which can lead to further gains.
Model Method Wiki2$\downarrow$ C4$\downarrow$ Task Avg.$\uparrow$
Mistral-7B-v0.3 Dense 4.82 7.72 68.7
Uniform 5.54 8.57 66.3
DP 5.79 8.84 66.0
EvoPress 5.21 8.42 67.1
Llama-3-8B Dense 5.54 7.10 68.6
Uniform 12.19 15.76 60.2
DP 29.00 20.03 61.3
EvoPress 7.49 12.03 64.3
Phi-3-Medium-14B Dense 4.02 8.31 73.2
Uniform 5.18 9.05 70.0
DP 5.72 9.71 69.1
EvoPress 5.09 9.00 70.8
@@ 4.3 Application 3: Quantization
Finally, we apply EvoPress to the more challenging task of non-uniform neural network quantization, where the widely adopted baseline is uniform per-layer quantization (Frantar et al., 2022; Lin et al., 2023; Chee et al., 2023). Additionally, we consider a DP-based approach for comparison. (While OWL has also been applied to quantization, the authors found that it underperforms even relative to uniform per-layer quantization (Yin et al., 2024).) The DP search is very similar to SPDY (Frantar & Alistarh, 2022), where the goal is to minimize the Normalized Mean Squared Error (NMSE), defined as $\mathrm{NMSE}=\|\hat{Y}-Y\|_{2}^{2}/\|Y\|_{2}^{2}$, where $Y$ represents the original model output at a layer, and $\hat{Y}$ the output of the compressed model. Then, the optimal compression allocation can be determined via a dynamic programming (DP) approach. The full SPDY method applies a second iterative random search step, which is very expensive to implement at LLM scale, and is therefore omitted.
@@
Figure 4: EvoPress enables non-uniform quantization with end-to-end speedups in vLLM. vLLM supports only 4-bit and 8-bit quantization, while requiring the query/key/value matrices as well as the up/gate projections to share the same bitwidth. The displayed average bitwidths include the overhead from groupwise scales.
@@
To validate the effectiveness of EvoPress, we consider the challenging problem of quantization to 3 bits and below. For this compression rate, uniform GPTQ quantization faces significant performance drops, motivating a dynamic quantization bitwidth allocation. We produce quantization levels at 2, 3, 4, 5, and 6 bits and search for an optimal compression profile with respect to the fitness function. The results in Table 3 indicate that non-uniform quantization with EvoPress produces superior models compared to the baseline methods. The improvements are even more pronounced at 2.25 bit and 2.5 bit quantization, as detailed in Appendix F.1. To show the real-world applicability of EvoPress, we additionally tested a restricted search limited to vLLM-supported options. vLLM allows only 4-bit and 8-bit quantization, and requires uniform bitwidth for query/key/values matrices and up/gate projections within a block. Figure 4 demonstrates strong mixed precision quantization even under these constraints.
We visualize the configurations found by EvoPress for Llama-3.1-8B in Appendix Figures 15 and 16. We observe that the second and final blocks are compressed less aggressively, while the first block undergoes the highest compression. This contrasts with unstructured sparsity, where the first block is among the least compressed. Therefore, dynamic compression allocation must account for the specific compression method used, which underscores the importance of automated compression allocation.
Overall, we observe that EvoPress yields significant accuracy improvements (e.g., 4.1 on the zero-shot averages for Llama-3-8B), compared to the uniform profile. Moreover, the improvement over the next-best method is always significant, both in terms of perplexity and zero-shot accuracy.
@@ 5 Conclusion
We have presented EvoPress, an optimization framework for non-uniform compression. EvoPress is based on a new evolutionary search algorithm with low sample and iteration complexity, especially well-suited to loss landscapes in LLM compression. Specifically, we have shown that EvoPress can converge extremely fast to accurate configurations for various non-uniform LLM compression problems, and is also fast to execute in practice. We also emphasize the breadth of our study, our method was implemented and tested on three different compression approaches, relative to prior work which largely focused on a single application. Interesting directions we did not investigate are 1) combining different compression approaches into the same search space, and 2) finer-grained structured pruning. We plan to investigate this in future work.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 A Warm-Up Argument for a Single Offspring
The overall goal of this section is to prove Theorem 3.1. As the main argument is quite complex, relying heavily on stochastic drift analysis, we begin with a warm-up, namely by presenting a simpler proof for the restricted case where $\lambda=1$.
@@
Theorem A.1 (Single offspring).
@@ Discussion.
It is important to note that the above proof relies on the key assumption that whenever one of the $j^{2}$ “good” pairs is selected during mutation, the resulting offspring is the fittest among all candidates. This condition holds naturally when there is only a single offspring, as the offspring produced by flipping one of the $j^{2}$ pairs will have higher fitness than the parent. However, in the case of multiple offspring, this approach breaks down, as an offspring produced by flipping one of the $j^{2}$ “good” pairs might still have lower fitness than another offspring that was not generated by flipping one of these $j^{2}$ “good” pairs.
@@ A.2 The Main Argument
Drift analysis, originally developed to study all kinds Markov chains, has become the most widely used technique for analyzing the runtime of evolutionary algorithms in recent years. It works by first defining a potential function $X^{(t)}$ that measures the progress over each step $t$ of the optimization. By estimating how this potential changes at each step in expectation, i.e., computing the drift in $X^{(t)}$, one can then make probabilistic statements about the number of steps required until the potential reaches a certain threshold, also called the hitting time. To this end, a variety of drift theorems have been established, two of which will be employed in our proof. For a more thorough introduction to drift analysis, we refer to Lengler (2020).
@@
Theorem A.2 (Multiplicative Drift, Tail Bound (Doerr & Goldberg, 2010)).
@@
Theorem A.3 (Variable Drift Theorem (Johannsen, 2010)).
@@
As in the proof of Theorem A.1 let $w\in\mathbb{R}^{n}$ represent the weights associated with a linear function of the form $f(x)=\sum_{i=1}^{n}x_{i}\cdot w_{i}$. To establish an upper bound, we can again assume that no two weights are equal. Additionally, without loss of generality, assume that the weights are ordered in increasing value, i.e., $w_{1}<w_{2}<\dots<w_{n}$, and that $k\leq n-k$, as the other case follows by symmetry. Let $x^{(t)}$ denote the search point at step $t$, and define
@@
which captures the number of inversions at step $t$. Since $x_{\text{opt}}=0^{k}1^{n-k}$ is the only bit string with $k$ 00-bits without inversions, we have $X^{(t)}=0$ if and only if $x^{(t)}=x_{\text{opt}}$. At the same time, no bit string with $k$ 00-bits has more than $k(n-k)$ inversions, hence, $X^{(t)}\leq k(n-k)$ at all times. During mutation, each of the $\lambda$ offspring is generated independently by copying the parent $x^{(t)}$, choosing uniformly at random one of the $1$-bits, choosing uniformly at random one of the 00-bits and finally flipping both bits. This flipping can also be viewed as switching both bits, so that bits “move” across the search point in consecutive generations. We will use this abstraction in a later step of the proof.
@@
At the same time, switching two bits corresponding to an inversion decreases the number of inversions by the difference in their positions, which we call the spread of an inversion. This implies that any offspring fitter than its parent must have fewer inversions than its parent and therefore, $X^{(t+1)}\leq X^{(t)}$ for all $t$. Note that we cannot make the same statement about the entire group of offspring, meaning, the fittest offspring is not guaranteed to have the fewest inversions. Since $X^{(t)}$ is non-increasing we can decompose $T$ into the number of steps required until for the first time the current search point $x^{(t)}$ has at most $k(n-k)/\lambda$ inversions and the number of steps required from there until the optimum is found. By linearity of expectation
@@
Denote by $\hat{T_{j}}$ the number of steps required until $Z_{j}$ reaches $1$, only counting steps where $Z_{j}$ is decreased. Using a tail bound for the Multiplicative Drift Theorem (Theorem A.2) we have that
@@
Finally, as the probability of having less than $Ck(n-k)/\lambda$ “successful” generations in the considered time period is sub-constant, we conclude via another union bound that there exists a constant $C^{\prime}$ such that event $B_{C}$ occurs with probability at least $1/C^{\prime}$. Consequently, we have
@@
Lemma A.4 indicates that when selecting an offspring uniformly at random from all those with higher fitness than the parent (i.e., those generated by flipping an inversion), the expected number of inversions in that offspring is at least $\sqrt{s}/16$ fewer than in the parent. We can now reveal the randomness in two steps. First, we only uncover how many of the generated offspring are fitter than the parent. Given that there is only a single fitter offspring, i.e., conditioned on $A_{1}$, we then uncover its number of inversions. Clearly, this single fitter offspring is now sampled uniformly at random from all offspring with higher fitness than $x^{(t)}$; thus, for $s\leq k(n-k)/\lambda$
@@
Finally, applying Johannsen’s Variable Drift Theorem (Johannsen, 2010) (Theorem A.3) yields